# Relaxed bound on performance of quantum key repeaters and secure content of generic private and independent bits

### 29 Feb 2024 v3

@@ III Technical preliminaries
We recall in this section certain facts and definitions and fix notation. A private state is a state of the form
$\gamma_{ABA^{\prime}B^{\prime}}=\frac{1}{d_{k}}\sum_{ij=0}^{d_{k}-1}|ii\rangle\langle jj|\otimes U_{i}\rho_{A^{\prime}B^{\prime}}U_{j}^{\dagger}\equiv$
$\tau|\psi_{+}\rangle\langle\psi_{+}|\otimes\rho_{A^{\prime}B^{\prime}}\tau^{\dagger}$ (9)
where $U_{i}$ are unitary transformations, and $\rho_{A^{\prime}B^{\prime}}$ is an arbitrary state on system $A^{\prime}B^{\prime}$ of dimension $d_{s}\otimes d_{s}$. A controlled unitary transformation $\tau:=\sum_{i}|ii\rangle\langle ii|\otimes U_{i}$ where $U_{i}$ are unitary transformations, is called a twisting and $|\psi_{+}\rangle\langle\psi_{+}|=\sum_{i,j=0}^{d_{k}-1}\frac{1}{d_{k}}|ii\rangle\langle jj|$. In the case of $d_{k}=2$, it can be obtained from Eq. (2) by tracing out system $E$. System $AB$ is called the key part, and system $A^{\prime}B^{\prime}$ is called the shield. We denote by $\gamma^{k}$ a private state with $2^{k}\otimes 2^{k}$ key part, i.e. containing at least $k$ key bits. In the case of $k=1$, the private state is called a private bit, or pbit, while for $k>1$, it is called a pdit (with $d_{k}=2^{k}$). A state $\rho$ is an $\epsilon$-approximate private state $\gamma_{d_{k}}$ if there exists $\epsilon>0$ such that $||\rho-\gamma_{d_{k}}||_{tr}\leq\epsilon$.
@@
where $\rho\approx_{\epsilon}\sigma$ denotes $||\rho-\sigma||_{tr}\leq\epsilon$, i.e. closeness is the trace-norm distance by $\epsilon$, with $||X||:=Tr\sqrt{XX^{\dagger}}$ and ${\cal P}$ is a completely positive trace-preserving map from the set of LOCC operations.
Technical, but important role in our considerations plays the so-called key attacked private state denoted as $\hat{\gamma}$. This is a private state that got measured on its key part and takes form
$\hat{\gamma}:=\frac{1}{d_{k}}\sum_{ij=0}^{d_{k}-1}|ii\rangle\langle ii|\otimes U_{i}\rho_{A^{\prime}B^{\prime}}U_{i}^{\dagger}.$ (11)
We note here, that $\hat{\gamma}$ is separable iff $U_{i}\rho U_{i}^{\dagger}$ are separable for $i\in\{0,\ldots,d_{k}-1\}$. We aim at generalizing results of Christandl and Ferrara (2017) to the case when we do not know if a certain key attacked state is separable.
Below we recall the main definitions and results from Christandl and Ferrara (2017). There the notion of key correlated states is introduced. this class of states is the one for which the bound for quantum key repeaters is given in Christandl and Ferrara (2017). Denoting by ${\cal Z}_{d_{k}}$ the Weyl operator of dimension $d_{k}$ of the form ${\cal Z}_{d_{k}}=\sum_{l=0}^{d_{k}-1}e^{i2\pi l/d_{k}}|l\rangle\langle l|$ we obtain the generalized Bell states as $\phi_{j}:={\cal Z}_{d_{k}}^{j}\otimes\mathds{1}_{B}(\phi)$ where $\phi:={1\over\sqrt{d_{k}}}\sum_{i}|i\rangle_{A}|i\rangle_{B}$. Then, the key correlated state takes form:
@@
The key attacked state of a key correlated state $\rho^{key-cor}_{ABA^{\prime}B^{\prime}}$ is the state obtained from it by applying local von-Neumann measurements on systems $A$ and $B$ in the computational basis. It takes the following form:
@@
One of the main results of Christandl and Ferrara (2017) connects the problem of distinguishability of a key correlated state from its key attacked version. It states that the one-way distillable entanglement of the key correlated state $\rho$ is quantified by means of the so-called locally measured-relative entropy ”distance” between the $\rho$ and its key attacked version when the latter state is separable. By locally measured relative entropy distance between two states we mean the relative entropy of two states measured both by a quantum completely positive, trace preserving map ${\cal M}$ (not necessarily a POVM as it was defined in Piani (2009)) on Alice’s side:
@@
where $D(\rho||\sigma)=-\mathrm{Tr}\rho\log\rho-{\mathrm{T}r}\rho\log\sigma$. In what follows, we need a regularized version of it, which reads:
@@
where in the above ${\cal M}$ acts naturally on all the $n$ subystems $A$ of $\rho$ (and $\sigma$ respectively).
@@
$E_{D}^{\rightarrow}(\rho)=D^{\infty}_{A}(\rho||\hat{\rho}).$ (18)
The central role that plays in our considerations is a variant of the key repeater rate. The one-way key repeater rate is defined in Christandl and Ferrara (2017), as follows, where the protocols consist of Local Operations and Classical Communication LOCC operations that are a composition of (i) local quantum operation on system $C_{1}C_{2}\equiv C$ followed by one-way classical communication, i.e. the one from $C_{1}C_{2}\equiv{\bf C}$ to $AB$ followed by (ii) arbitrary LOCC operation between $A$ and $B$. We denote this composition as $LOCC(C_{1}C_{2}\rightarrow A\leftrightarrow B)$.
@@
(cf. Appendix G in Christandl and Ferrara (2017)) Consider any two bipartite states $\rho_{AC_{1}}$ of systems $AC_{1}$ and $\rho^{\prime}_{C_{2}B}$ of systems $C_{2}B$ (not necesarily equal), called further input states. A protocol ${\cal P}$ consisting of $LOCC(C_{1}C_{2}\rightarrow A\leftrightarrow B)$ operations which transform $n$ copies of the input states into $\epsilon$ approximation (in trance norm distance) of a private state $\gamma^{k}$ shared between $A$ and $B$ has the key rate $\frac{k}{n}$. The maximal value of the key rates over such protocols in the asymptotic limit of large $n$ and vanishing error $\epsilon$ is is called a one-way key repeater rate of $\rho_{AC_{1}}\otimes\rho_{C_{2}B}^{\prime}$, and reads
@@
Following Christandl and Ferrara (2017) we will use the locally measured on system ${\bf C}\equiv C_{1}C_{2}$, regularized relative entropy applied to the scenario of key repeaters:
$D_{\bf{C}}^{\infty}(\rho_{AC_{1}}\otimes\rho^{\prime}_{C_{2}B}||\rho^{\prime\prime})\coloneqq\lim_{n\rightarrow\infty}{1\over n}$
@@
which is nothing but the locally measured relative entropy with $A$ in Eq. (15) identified with the system ${\bf C}=C_{1}C_{2}$ and $B$ with $AB$ respectively, for the input state $\rho_{AC_{1}}\otimes\rho_{C_{2}B}$.Let us emphasise here, that ${\cal M}_{\bf C}$ acts locally on all $n$ subsystems $C_{1}C_{2}$ of the state $(\rho_{AC_{1}}\otimes\rho_{C_{2}B})^{\otimes n}$ and analogously on all $n$ subsystems $C_{1}C_{2}$ of the state ${\rho^{\prime\prime}_{AC_{1}C_{2}B}}^{\otimes n}$.
@@
for any $\sigma$ separable in $ACC_{1}:B$ or $A:CC_{2}B$ cut. Then for $\rho,\rho^{\prime}$ being key correlated states, and by choosing $\sigma=\hat{\rho}\otimes\hat{\rho^{\prime}}$ (i.e. assuming that the key attacked the correlated state of at least one of the two states is separable) in Christandl and Ferrara (2017) they have obtained from Theorem 18 invoked above
$R^{C_{1}C_{2}\rightarrow AB}(\rho_{AC_{1}},\rho^{\prime}_{C_{2}B})\leq E_{D}^{C_{1}\rightarrow A}(\rho_{AC_{1}}\otimes\rho^{\prime}_{C_{2}B})$ (22)
@@
[…]
@@
In Hayashi (2017), based on the above quantity, the sandwiched relative entropy is defined as follows:
@@
As the second part of this work heavily utilizes random quantum objects, we provide a crash course on random matrices in Appendix B.
Finally, we would like to recall the results of Yang et al. (2019) on private randomness distillation, as our result will also apply to this resource. There the notion of independent states was introduced. These states take the form of twisted coherence, just like the private state is twisted entanglement.
@@
$\alpha_{d_{A},d_{B}}:={1\over d_{A}d_{B}}\sum_{i,j=0}^{d_{A}-1}\sum_{k,l=0}^{d_{B}-1}|i\rangle\langle j|_{A}\otimes|k\rangle\langle l|_{B}\otimes U_{ik}\sigma_{A^{\prime}B^{\prime}}U_{jl}^{\dagger}$ (27)
where the unitary transformation $\tau=\sum_{ij}|ij\rangle\langle ij|\otimes U_{ij}$ is the twisting, and $\sigma_{A^{\prime}B^{\prime}}$ is an arbitrary state on $A^{\prime}B^{\prime}$ system.
The independent state has the property that measured on system $AB$ yields $\log_{2}d_{A}$ bits of ideally private randomness for Alice and $\log_{2}d_{B}$ for Bob. The randomness is ideally private since the construction of the state assures that the outcomes of measurements on the key part system $AB$ are decorrelated from the purifying system $E$. We also recall the scenario of distributed randomness distillation of Yang et al. (2019). In the latter scenario, two parties are trying to distil locally private randomness, which is independent for each party and decoupled from the system of the environment. The honest parties share a dephasing channel via which they can communicate. One can also consider the case in which the honest parties have local access to an unlimited number of maximally mixed states. The parties distil private randomness in the form of independent states. We recall below the main results of Yang et al. (2019):
@@
    for no communication and no noise, $R_{A}\leq\log|A|-S(A|B)_{+}$, $R_{B}\leq\log|B|-S(B|A)_{+}$, and $R_{A}+R_{B}\leq R_{G}$, where $[t]_{+}=\max\{0,t\}$;
@@
    for free communication but no noise, $R_{A}\leq\log|AB|-\max\{S(B),S(AB)\}$, $R_{B}\leq\log|AB|-\max\{S(A),S(AB)\}$, and $R_{A}+R_{B}\leq R_{G}$.
Further, the rate regions in settings 1), 2), 3) are tight.
@@
is the quantum mutual information. We will sometimes neglect the subscript $XY$ in the above if understood from the context. Recently, it has been shown that private states are independent states Horodecki et al. (2020). However, naturally, the set of independent states is strictly larger. In particular, a local independent bit is not necessarily a private state:
$\alpha_{d_{A}}:=\tau|+\rangle\langle+|\otimes\sigma_{A^{\prime}B^{\prime}}\tau^{\dagger},$ (29)
@@ IV Relaxed bound on one-way private key repeaters
In this section, we provide relaxed bound on one-way private key repeaters. It holds for the so-called key correlated states $\rho_{key}$ Christandl and Ferrara (2017). The bound takes the form of the regularized, measured relative entropy of entanglement from any state scaled by the factor $\alpha\over\alpha-1$ increased by the $\alpha$-sandwiched relative entropy of entanglement $E_{\alpha}$ of the key-attacked version of $\rho_{key}$. Although the latter bound can not be tighter than the bound by relative entropy of entanglement Horodecki et al. (2005, 2009a), it has appealing form, as it in a sense, computes the distance from a separable state via a ’proxy’ state, which can be arbitrary.
@@
As the main technical contribution, we will first present the lemma, which upper bounds the fidelity with a singlet of a state which up to its inner structure (being outcome of certain protocol) is arbitrary. Such a bound was known so far for the fidelity of singlet with the so-called twisted separable states (see lemma 7 of Horodecki et al. (2009a)) and proved useful in providing upper bounds on distillable key (see Khatri and Wilde (2020) and references therein). Our relaxation is based on the strong-converse bound for the private key, formulated in terms of the $E_{\alpha}$.
@@
We begin by explaining the idea of a technical lemma which we present below. It is known that any $d\otimes d$ separable state is bounded away as $1\over d$ from a maximally entangled state of local dimension $d$ in terms of fidelity Horodecki et al. (2009a). However, an arbitrary state can not be bounded arbitrarily away from the singlet state in terms of fidelity. There must be a penalty term that shows how fast this fidelity grows when the state is increasingly entangled, i.e., far and far from being separable. Such a term follows from the strong converse bound on the distillable key. Indeed, this is the essence of $\tilde{E}_{\alpha}$ being the strong converse bound: If one tries to distil more key than the $\tilde{E}_{\alpha}$ from a state, then the fidelity with singlet of the output state increases exponentially fast to $1$, in terms of $\tilde{E}_{\alpha}$. In our case, the state compared with the singlet in terms of fidelity will be the state after the performance of the key repeater protocol subjected to twisting and partial trace. Namely, a state $\sigma^{\otimes n}$ of systems $ABC$ after the action of a map $\Lambda\in LOCC({\bf C}\rightarrow(A\leftrightarrow B))$ which can be now of systems $ABA^{\prime}B^{\prime}C$, traced over $C$ and further rotated by a twisting unitary transformation $\tau$ and traced over system $A^{\prime}B^{\prime}$. The last two actions, i.e., rotation and twisting, form just a mathematical tool that allows checking if the state under consideration, here the protocol’s output, is close to a private state. Indeed, a private state, after performing the inverse operation to twisting and the partial trace becomes the singlet state Horodecki et al. (2005). We note here, that the proof technique in the Lemma (1) below, bare partial analogy to the proof of the fact that the so called second type distillable entanglement is upper bounded by the relative entropy of entanglement as shown in Vedral and Plenio (1998) and described in Theorem $8.7$ of Hayashi (2017).
@@
In what follows, we directly use the proof of the strong-converse bound for the distillable key of Das et al. (2021), which is based on Wilde et al. (2017); Das et al. (2020); Christandl and Müller-Hermes (2017). Let us suppose $\operatorname{Tr}\tilde{\sigma}\psi_{+}^{nR}=1-\epsilon$. Let us also choose a state $\hat{\sigma}$, to be a twisted separable state of the form $\hat{\sigma}=\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\operatorname{Tr}_{C}\Lambda(\sigma_{A{\bf C}:B}^{\prime})\tau^{\dagger}$ with $\sigma^{\prime}$ being arbitrary separable state in $(A{\bf C}:B)$ cut. Note here that by definition $\Lambda$ preserves separability in $A{\bf C}:B$ cut, and hence in $A:B$ as well.
Since any such state has overlap with the singlet state $\psi^{nR}_{+}$ less than $\frac{1}{2^{nR}}\equiv 1/K$ (see lemma 7 of Horodecki et al. (2009a)), we have by taking ${\mathrm{P}}=|\psi_{+}^{nR}\rangle\langle\psi_{+}^{nR}|$ definition of $D^{\epsilon}_{h}$ given in Eq. (23) a bound:
$\log_{2}K\leq D^{\epsilon}_{h}(\tilde{\sigma}||\hat{\sigma}).$ (31)
We then upper bound $D^{\epsilon}_{h}$ by $\tilde{D}_{\alpha}$ as follows, for every $\alpha\in(1,\infty)$ (see A8 in the Appendix of Das et al. (2021)):
@@
[…]
@@
We can further drop also the operation $\Lambda$ again using monotonicity of the relative Renyi divergence $\tilde{D}_{\alpha}$. We thus arrive at
@@
Since $\sigma^{\prime}$ is an arbitrary separable state, we can take also infimum over this set obtaining:
@@
$1-\epsilon\leq 2^{-n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}E_{\alpha}(\sigma^{\otimes n}))}.$ (36)
@@
For a maximally entangled state $\psi_{+}^{Rn}$, a state $\tilde{\sigma}_{n}:=\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\operatorname{Tr}_{C}\Lambda(\sigma^{\otimes n})\tau^{\dagger}$ where $\tau$ is a twisting, for every $\alpha\in(1,\infty)$, there is
[…]
$n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}E_{\alpha}(\sigma^{\otimes n})).$ (38)

### 09 Mar 2026 v4

@@ III Technical preliminaries
In this section, we recall certain facts and definitions and fix notation. A private state is a state of the form
$\gamma_{ABA^{\prime}B^{\prime}}=\frac{1}{d_{k}}\sum_{ij=0}^{d_{k}-1}|ii\rangle\langle jj|_{AB}\otimes U_{i}\rho_{A^{\prime}B^{\prime}}U_{j}^{\dagger}\equiv$
$\tau|\psi_{+}\rangle\langle\psi_{+}|_{AB}\otimes\rho_{A^{\prime}B^{\prime}}\tau^{\dagger}$ (10)
where $U_{i}$ are unitary transformations, and $\rho_{A^{\prime}B^{\prime}}$ is an arbitrary state on system $A^{\prime}B^{\prime}$ of dimension $d_{s}\otimes d_{s}$. A controlled unitary transformation $\tau:=\sum_{i}|ii\rangle\langle ii|\otimes U_{i}+\sum_{i\neq j}|ij\rangle\langle ij|\otimes\mathds{1}$ where $U_{i}$ are unitary transformations, is called a twisting and $|\psi_{+}\rangle\langle\psi_{+}|=\sum_{i,j=0}^{d_{k}-1}\frac{1}{d_{k}}|ii\rangle\langle jj|$. In the case of $d_{k}=2$, it can be obtained from Eq. (2) by tracing out system $E$. System $AB$ is called the key part, and system $A^{\prime}B^{\prime}$ is called the shield. We denote by $\gamma^{k}$ a private state with $2^{k}\otimes 2^{k}$ key part, i.e., containing at least $k$ key bits. In the case of $k=1$, the private state is called a private bit, or pbit, while for $k>1$, it is called a pdit (with $d_{k}=2^{k}$). A state $\rho$ is an $\epsilon$-approximate private state $\gamma_{d_{k}}$ if there exists $\epsilon>0$ such that $\rho-\gamma_{d_{k}}||_{tr}\leq\epsilon$.
@@
where $\rho\approx_{\epsilon}\sigma$ denotes $\rho-\sigma||_{tr}\leq\epsilon$, i.e. closeness is the trace-norm distance by $\epsilon$, with $X||:=Tr\sqrt{XX^{\dagger}}$ and $\cal P$} is a completely positive trace-preserving map from the set of LOCC operations.
A technical, but important, role in our considerations is played by the so-called key attacked private state, denoted as $\hat{\gamma}$. This is a private state that got measured on its key part and takes the form
$\hat{\gamma}:=\frac{1}{d_{k}}\sum_{i=0}^{d_{k}-1}|ii\rangle\langle ii|_{AB}\otimes U_{i}\rho_{A^{\prime}B^{\prime}}U_{i}^{\dagger}.$ (12)
We note here, that $\hat{\gamma}$ is separable iff $U_{i}\rho U_{i}^{\dagger}$ are separable for $i\in\{0,\ldots,d_{k}-1\}$. We aim at generalizing the results of Christandl and Ferrara (2017) to the case when we do not know if a certain key attacked state is separable.
Below, we recall the main definitions and results from Christandl and Ferrara (2017). There, the notion of key correlated states is introduced. This class of states is the one for which the bound for quantum key repeaters is given in Christandl and Ferrara (2017). Denoting by ${\cal Z}_{d_{k}}$ the Weyl operator of dimension $d_{k}$ of the form ${\cal Z}_{d_{k}}=\sum_{l=0}^{d_{k}-1}e^{i2\pi l/d_{k}}|l\rangle\langle l|$ we obtain the generalized Bell states as $\phi_{j}:={\cal Z}_{d_{k}}^{j}\otimes\mathds{1}_{B}(\phi)$ where $\phi:={1\over\sqrt{d_{k}}}\sum_{i}|i\rangle_{A}|i\rangle_{B}$. Then, the key correlated state takes form:
@@
The key attacked state $\hat{\rho}_{ABA^{\prime}B^{\prime}}$ of a key correlated state $\rho^{key-cor}_{ABA^{\prime}B^{\prime}}$ defined in equation (13) is the state obtained from $\rho^{key-cor}_{ABA^{\prime}B^{\prime}}$ by applying local von-Neumann measurements on systems $A$ and $B$ in the computational basis. It takes the following form:
@@
[…]
@@
where $D(\rho||\sigma)=\mathrm{Tr}\rho\log\rho-{\mathrm{T}r}\rho\log\sigma$. In what follows, we need a regularized version of it, which reads:
@@
[…]
@@
[…]
A specific variant of the key repeater rate plays a central role in our considerations. The one-way key repeater rate is defined in Christandl and Ferrara (2017), as follows, where the protocols consist of Local Operations and Classical Communication LOCC operations that are a composition of (i) local quantum operation on system $C_{1}C_{2}\equiv C$ followed by one-way classical communication, i.e., the one from $C_{1}C_{2}\equiv{\bf C}$ to $AB$ followed by (ii) an arbitrary LOCC operation between $A$ and $B$. We denote this composition as $LOCC(C_{1}C_{2}\rightarrow A\leftrightarrow B)$.
@@
(cf. Appendix G in Christandl and Ferrara (2017)) Consider any two bipartite states $\rho_{AC_{1}}$ of systems $AC_{1}$ and $\rho^{\prime}_{C_{2}B}$ of systems $C_{2}B$ (not necessarily equal), called further input states. A protocol $\cal P$} consisting of $LOCC(C_{1}C_{2}\rightarrow A\leftrightarrow B)$ operations which transform $n$ copies of the input states into $\epsilon$ approximation (in trace norm distance) of a private state $\gamma^{k}$ shared between $A$ and $B$ has the key rate $\frac{k}{n}$. The maximal value of the key rates over such protocols in the asymptotic limit of large $n$ and vanishing error $\epsilon$ is called a one-way key repeater rate of $\rho_{AC_{1}}\otimes\rho_{C_{2}B}^{\prime}$, and reads
@@
Following Christandl and Ferrara (2017), we will use the locally measured on system ${\bf C}\equiv C_{1}C_{2}$, regularized relative entropy applied to the scenario of key repeaters:
$D_{\bf{C}}^{\infty}(\rho_{AC_{1}}\otimes\rho^{\prime}_{C_{2}B}||\rho_{AC_{1}C_{2}B}^{\prime\prime})\coloneqq\lim_{n\rightarrow\infty}{1\over n}$
@@
which is nothing but the locally measured relative entropy with $A$ in Eq. (16) identified with the system ${\bf C}=C_{1}C_{2}$ and $B$ with $AB$ respectively, for the input state $\rho_{AC_{1}}\otimes\rho_{C_{2}B}$. Let us emphasize here, that ${\cal M}_{\bf C}$ acts locally on all $n$ subsystems $C_{1}C_{2}$ of the state $(\rho_{AC_{1}}\otimes\rho_{C_{2}B})^{\otimes n}$ and analogously on all $n$ subsystems $C_{1}C_{2}$ of the state ${\rho^{\prime\prime}_{AC_{1}C_{2}B}}^{\otimes n}$.
@@
for any $\sigma$ separable in $AC_{1}C_{2}:B$ or $A:C_{1}C_{2}B$ cut. Then for $\rho,\rho^{\prime}$ being key correlated states, and by choosing $\sigma=\hat{\rho}\otimes\hat{\rho^{\prime}}$ (i.e. assuming that the key attacked the correlated state of at least one of the two states is separable) in Christandl and Ferrara (2017) they have obtained from Theorem 19 invoked above
$R^{C_{1}C_{2}\rightarrow AB}(\rho_{AC_{1}},\rho^{\prime}_{C_{2}B})\leq E_{D}^{C_{1}C_{2}\rightarrow AB}(\rho_{AC_{1}}\otimes\rho^{\prime}_{C_{2}B})$ (23)
@@
and the sandwiched Rényi relative entropy Wilde et al. (2014); Müller-Lennert et al. (2013):
@@
In Hayashi (2017), based on the above quantity, the sandwiched Rényi relative entropy of entanglement is defined as follows:
@@
As the second part of this work heavily relies on random quantum objects, we provide a brief introduction to random matrices in Appendix A.
Finally, we would like to recall the results of Yang et al. (2019) on private randomness distillation, as our result also applies to this resource. There, the notion of independent states was introduced. These states take the form of twisted coherence, just like the private state is twisted entanglement.
@@
$\alpha_{d_{A},d_{B}}:=$ (28)
$\qquad{1\over d_{A}d_{B}}\sum_{i,j=0}^{d_{A}-1}\sum_{k,l=0}^{d_{B}-1}|i\rangle\langle j|_{A}\otimes|k\rangle\langle l|_{B}\otimes U_{ik}\sigma_{A^{\prime}B^{\prime}}U_{jl}^{\dagger}\equiv$
$\qquad\qquad\tau|+_{d_{A}}\rangle\langle+_{d_{B}}|_{A}\otimes|+_{d_{B}}\rangle\langle+_{d_{B}}|_{B}\otimes\sigma_{A^{\prime}B^{\prime}}\tau^{\dagger}$
where the unitary transformation $\tau=\sum_{ij}|ij\rangle\langle ij|\otimes U_{ij}$ is the twisting, $\sigma_{A^{\prime}B^{\prime}}$ is an arbitrary state on $A^{\prime}B^{\prime}$ system and $|+_{d}\rangle=1/\sqrt{d}\sum_{i=0}^{d-1}|i\rangle$.
The independent state has the property that, when measured on the system $AB$, yields $\log_{2}d_{A}$ bits of ideally private randomness for Alice and $\log_{2}d_{B}$ for Bob. The randomness is ideally private since the construction of the state assures that the outcomes of measurements on the key part system $AB$ are decorrelated from the purifying system $E$. We also recall the scenario of distributed randomness distillation of Yang et al. (2019). In the latter scenario, two parties are trying to distill locally private randomness, which is independent for each party and decoupled from the environment’s system. The honest parties share a dephasing channel via which they can communicate (free communication). One can also consider the case in which the honest parties have local access to an unlimited number of maximally mixed states (free noise). The parties, $A$ and $B$, distill private randomness at rates $R_{A}$ and $R_{B}$, respectively, in the form of independent states. We recall below the main results of Yang et al. (2019):
@@
    for no communication and no noise, $R_{A}\leq\log|A|-S(A|B)_{+}$, $R_{B}\leq\log|B|-S(B|A)_{+}$, and $R_{A}+R_{B}\leq R_{G}$, where $t]_{+}=\max\{0,t\}$;
@@
    for free communication but no noise, $R_{A}\leq\log|AB|-\max\{S(B),S(AB)\}$, $R_{B}\leq\log|AB|-\max\{S(A),S(AB)\}$, and $R_{A}+R_{B}\leq R_{G}$,
where $R_{G}$ denotes global purity and equals $\log|AB|-S(AB)$. Further, the rate regions in settings 1), 2), 3) are tight.
@@
is the quantum mutual information. We will sometimes neglect the subscript $XY$ in the above if it is understood from the context. Recently, it has been shown that private states are independent states Horodecki et al. (2020). However, naturally, the set of independent states is strictly larger. In particular, a local independent bit is not necessarily a private state:
$\alpha_{d_{A}}:=\tau|+\rangle\langle+|_{A}\otimes\sigma_{A^{\prime}B^{\prime}}\tau^{\dagger},$ (30)
@@ IV Relaxed bound on one-way private key repeaters
In this section, we provide a relaxed bound on one-way private key repeaters. It holds for the so-called key correlated states $\rho_{key}$ Christandl and Ferrara (2017). The bound takes the form of the regularized, measured relative entropy of entanglement from any state scaled by the factor $\alpha\over\alpha-1$ increased by the $\alpha$-sandwiched relative entropy of entanglement $E_{\alpha}$ of the key-attacked version of $\rho_{key}$. Although the latter bound can not be tighter than the bound by relative entropy of entanglement Horodecki et al. (2005, 2009a), it has appealing form, as it, in a sense, computes the distance from a separable state via a ’proxy’ state, which can be arbitrary.
@@
As the main technical contribution, we will first present the lemma, which upper bounds the fidelity with a singlet of a state which up to its inner structure (being outcome of certain protocol) is arbitrary. Such a bound was known so far for the fidelity of a singlet with the so-called twisted separable states (see Lemma 7 of Horodecki et al. (2009a)) and proved useful in providing upper bounds on distillable key (see Khatri and Wilde (2020) and references therein). Our relaxation is based on the strong-converse bound for the private key, formulated in terms of the $\tilde{E}_{\alpha}$.
@@
We begin by explaining the idea of a technical lemma, which we present below. It is known that any $d\otimes d$ separable state is bounded away as $1\over d$ from a maximally entangled state of local dimension $d$ in terms of fidelity Horodecki et al. (2009a). However, an arbitrary state can not be bounded arbitrarily away from the singlet state in terms of fidelity. There must be a penalty term that quantifies how rapidly this fidelity increases as the state becomes increasingly entangled, i.e., far and far from being separable. Such a term follows from the strong converse bound on the distillable key. Indeed, this is the essence of $\tilde{E}_{\alpha}$ being the strong converse bound: If one tries to distill more key than the $\tilde{E}_{\alpha}$ from a state, then the fidelity with a singlet of the output state increases exponentially fast to $1$, in terms of $\tilde{E}_{\alpha}$. In our case, the state compared with the singlet in terms of fidelity will be the state after the performance of the key repeater protocol, subjected to twisting and partial trace. Namely, a state $\sigma^{\otimes n}$ of systems $ABC$ after the action of a map $\Lambda\in LOCC({\bf C}\rightarrow(A\leftrightarrow B))$ which can be now of systems $ABA^{\prime}B^{\prime}C$, traced over $C$ and further rotated by a twisting unitary transformation $\tau$ and traced over system $A^{\prime}B^{\prime}$. The last two actions, i.e., rotation and twisting, serve only as a mathematical tool to check whether the state under consideration, here the protocol’s output, is close to a private state. Indeed, a private state, after performing the inverse operation to twisting and the partial trace, becomes the singlet state Horodecki et al. (2005). We note here that the proof technique in the Lemma (1) below bears a partial analogy to the proof of the fact that the so called second type distillable entanglement is upper bounded by the relative entropy of entanglement as shown in Vedral and Plenio (1998) and described in Theorem $8.7$ of Hayashi (2017).
@@
In what follows, we directly use the proof of the strong-converse bound for the distillable key of Das et al. (2021b), which is based on Wilde et al. (2017); Das et al. (2020); Christandl and Müller-Hermes (2017). Let us suppose $\operatorname{Tr}\tilde{\sigma}\psi_{+}^{nR}=1-\epsilon$. Let us also choose a state $\hat{\sigma}$, to be a twisted separable state of the form $\hat{\sigma}=\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\operatorname{Tr}_{C}\Lambda(\sigma_{A{\bf C}:B}^{\prime})\tau^{\dagger}$ with $\sigma^{\prime}$ being arbitrary separable state in $(A{\bf C}:B)$ cut. Note here that by definition $\Lambda$ preserves separability in $A{\bf C}:B$ cut, and hence in $A:B$ as well.
Since any such state has an overlap with the singlet state $\psi^{nR}_{+}$ less than $\frac{1}{2^{nR}}\equiv 1/K$ (see lemma 7 of Horodecki et al. (2009a)), we have by taking ${\mathrm{P}}=|\psi_{+}^{nR}\rangle\langle\psi_{+}^{nR}|$ definition of $D^{\epsilon}_{h}$ given in Eq. (24) a bound:
[…]
We then upper bound $D^{\epsilon}_{h}$ by $\tilde{D}_{\alpha}$ as follows, for every $\alpha\in(1,\infty)$ (see A8 in the Appendix of Das et al. (2021b)):
@@
We can now relax $\tilde{\sigma}$ to $\Lambda(\sigma^{\otimes n})$ due to monotonicity of the sandwich Rényi relative entropy distance under a jointly applied channel, so that
@@
We can also further drop the operation $\Lambda$ again using monotonicity of the relative Rényi divergence $\tilde{D}_{\alpha}$. We thus arrive at
@@
Since $\sigma^{\prime}$ is an arbitrary separable state, we can also take the infimum over this set, obtaining:
@@
$1-\epsilon\leq 2^{-n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}\tilde{E}_{\alpha}(\sigma^{\otimes n}))}.$ (37)
@@
Let $\sigma_{AA^{\prime}BB^{\prime}C}$ be an arbitrary quantum state. For a maximally entangled state $\psi_{+}^{Rn}$, a state $\tilde{\sigma}_{n}:=\operatorname{Tr}_{A^{\prime}B^{\prime}}\tau\operatorname{Tr}_{C}\Lambda(\sigma^{\otimes n})\tau^{\dagger}$ where $\tau$ is a twisting, and every $\alpha\in(1,\infty)$, there is
$E_{R}^{\mbox{Cone}(\tilde{\sigma}_{n})}(\psi_{+}^{Rn})\equiv\inf_{\sigma_{n}\in\mbox{Cone}(SEP,\tilde{\sigma}_{n})}D(\psi_{+}^{Rn}||\sigma_{n})\geq$
$n(\frac{\alpha-1}{\alpha})(R-\frac{1}{n}\tilde{E}_{\alpha}(\sigma^{\otimes n})).$ (39)